METAL-CONTAINING MOLECULES BEYOND THE SOLAR SYSTEM: A LABORATORY AND RADIO ASTRONOMICAL PERSPECTIVE

\underline{L. M. ZIURYS}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

Although the history of interstellar molecules began around 1970, with the millimeter-wave detection of CO in the Orion Nebula, metal-containing species have been somewhat elusive for astronomical searches. Only in the past two decades have metal-bearing molecules been identified in space, starting with metal halides (NaCl, KCl, AlCl, and AlF), and then metal cyanide and isocyanide species (MgNC, MgCN, NaCN, and AlNC). Moreover, the metal-containing molecules seemed to be present in a single astronomical object: the envelope of a dying, carbon-rich star, IRC+10216. However, with improvements both in laboratory spectroscopy and telescope sensitivity, it is becoming clear that the relevance of metalcontaining species in astrophysics is increasing. Metal oxide and hydroxide species, such as AlO and AlOH, have recently been identified in interstellar space. Metal-containing molecules are now being found in other astronomical sources, such as the oxygen-rich shell surrounding VY Canis Majoris, a supergiant star. These new astronomical discoveries will be presented, as well as the laboratory measurements that made them possible. New directions in rotational spectroscopy of metal-bearing molecules will also be discussed.